### v1

# SpeechJudge: Towards Human-Level Judgment for Speech Naturalness

Abstract
Aligning large generative models with human feedback is a critical challenge. In speech synthesis, this is particularly pronounced due to the lack of a large-scale human preference dataset, which hinders the development of models that truly align with human perception. To address this, we introduce SpeechJudge, a comprehensive suite comprising a dataset, a benchmark, and a reward model centered on naturalness—one of the most fundamental subjective metrics for speech synthesis. First, we present SpeechJudge-Data, a large-scale human feedback corpus of 99K speech pairs. The dataset is constructed using a diverse set of advanced zero-shot text-to-speech (TTS) models across diverse speech styles and multiple languages, with human annotations for both intelligibility and naturalness preference. From this, we establish SpeechJudge-Eval, a challenging benchmark for speech naturalness judgment. Our evaluation reveals that existing metrics and AudioLLMs struggle with this task; the leading model, Gemini-2.5-Flash, achieves less than 70% agreement with human judgment, highlighting a significant gap for improvement. To bridge this gap, we develop SpeechJudge-GRM, a generative reward model (GRM) based on Qwen2.5-Omni-7B. It is trained on SpeechJudge-Data via a two-stage post-training process: Supervised Fine-Tuning (SFT) with Chain-of-Thought rationales followed by Reinforcement Learning (RL) with GRPO on challenging cases. On the SpeechJudge-Eval benchmark, the proposed SpeechJudge-GRM demonstrates superior performance, achieving 77.2% accuracy (and 79.4% after inference-time scaling @10) compared to a classic Bradley-Terry reward model (72.7%). Furthermore, SpeechJudge-GRM can be also employed as a reward function during the post-training of speech generation models to facilitate their alignment with human preferences.
[…]
1 Introduction
The collection and integration of human feedback corpora for model alignment has become a critical stage in the development of modern large-scale generative models, proving indispensable in domains such as text [44, 37, 3], image [54, 30], and video generation [55, 33].
11footnotetext: Equal Contribution.
In the field of speech synthesis, naturalness has long been a cornerstone subjective metric for quality assessment [24, 2, 13, 56, 28], representing one of the most general-purpose indicators of performance [47, 45]. Prior research has explored automated speech assessment through MOS predictors [41, 21] and constructed the human feedback corpora for specific attributes like the low-level acoustic quality [50]. However, a large-scale human feedback corpus centered on the holistic quality of naturalness—and a corresponding reward model trained to capture these preferences—remains a notably underexplored area. To fill this void, this paper focuses on the dimension of speech naturalness and present a three-part contribution:
1. A Large-scale Human Feedback Dataset: SpeechJudge-Data. We recruit human annotators to provide feedback on synthesized speeches, with a focus on assessing two fundamental speech aspects: intelligibility and naturalness. For data synthesis, we employ a diverse set of advanced, open-source zero-shot TTS models with varying architectures (such as CosyVoice2 [14], Ints [61], F5-TTS [7], and MaskGCT [52]) to produce the compared speech pairs. We prepare speech references in both regular and expressive styles, construct multilingual target texts, and cover both monolingual and cross-lingual synthesis scenarios to ensure data diversity (Section 3.1). We instruct human annotators to perform two tasks based on a speech pair (Figure 1): (a) pointwise annotation of text accuracy to assess intelligibility, and (b) pairwise preference annotation to judge relative speech naturalness. This extensive effort, involving 69 labelers over two months, results in 99K annotated pairs, with each pair receiving an average of 2.49 annotations from different labelers. We believe the SpeechJudge-Data can serve as a valuable corpus for alignment research in speech synthesis (e.g., DPO alignment [39] or reward modeling [44, 37, 3] in Section 5).
2. An Evaluation Benchmark for Speech Naturalness Judgment: SpeechJudge-Eval. We design a dedicated evaluation benchmark for the task of speech naturalness judgment. The task is structured as follows: given a target text and two corresponding speech samples, a model needs to judge which one is more natural. To construct the evaluation set, we select a subset from the SpeechJudge-Data where human annotators demonstrated high inter-annotator agreement, ensuring a high-quality ground truth. We assess the naturalness judgment capabilities of a wide range of metrics and models, including Word Error Rate (WER) [38, 16], Fréchet Audio Distance (FAD) [27], MOS predictors [41, 40, 48], Deepfake Detectors [25, 49], and AudioLLMs [56, 28, 53, 8, 22]. Our evaluations reveal that even the most capable model—specifically, Gemini-2.5-Flash [8] in our experiments—achieved less than 70% agreement with human preferences. This finding highlights a significant performance gap and underscores the substantial room for research and improvement in automated speech naturalness judgment.
3. A Generative Reward Model for Speech Naturalness: SpeechJudge-GRM. To develop a reward model that more effectively captures human preferences, we develop SpeechJudge-GRM, a generative reward model (GRM) [59, 34] trained on the SpeechJudge-Data. Specifically, we base our model on Qwen2.5-Omni-7B [56] and design a two-stage post-training process. During the first stage, we perform SFT as the “cold start” to improve the model’s instruction-following and rationale-based reasoning capabilities. To achieve this, we leverage Gemini-2.5-Flash [8] to generate Chain-of-Thought (CoT) data for speech naturalness judgment task. In the second stage, we focus on more challenging cases of SpeechJudge-Data, which we define as instances where Gemini-2.5-Flash fails to make the correct judgment. Treating the human-annotated labels as the verifiable reward [11, 34], we apply the GRPO-based RL stage [42]. Our experiments demonstrate that when trained on the same data, SpeechJudge-GRM significantly outperformed the classic Bradley-Terry reward model (BTRM) [4, 39], achieving a higher accuracy in predicting human preferences (77.2% for SpeechJudge-GRM vs. 72.7% for SpeechJudge-BTRM, Table 3). Besides, SpeechJudge-GRM also supports inference-time scaling and offers explainability through its CoT outputs. Furthermore, SpeechJudge-GRM can also be employed as an objective naturalness metric for sample selection (Figure 5) or as a reward function in RL algorithms to enhance the quality of existing speech generation models (Figure 6).
2 Related Work
Human Alignment for Speech Generation Aligning generative models with human feedback has proven crucial, a process also known as RLHF in LLMs [37, 3]. In the vision domain, many similar human preference datasets exist, such as Pick-a-Pic [30], ImageReward [54], and VideoReward [33]. The speech synthesis field, pioneering efforts to construct human corpora involved MOS datasets [41, 21]. However, these datasets often did not use advanced TTS models for data generation, provided only the pointwise labels rather than the direct pairwise human preference, and were limited in scale. More recently, efforts have focused on building human feedback corpora centered on specific speech attributes, such as low-level acoustic quality [50], intelligibility [61], or the instruction-following capabilities of spoken dialogue systems [23, 17]. Despite this progress, a large-scale human feedback corpus built specifically around naturalness—one of the most general-purpose and fundamental metrics for speech synthesis [47, 45]—has remained a critical missing piece.
AudioLLM as a Judge Using LLMs as automated quality evaluators is a prominent topic in the textual LLM field, popularized by the “LLM-as-a-judge” paradigm [64]. This approach has recently been extended to the speech domain. A concurrent work, AudioJudge [36], evaluates the capabilities and limitations of using AudioLLMs for speech quality assessment and paralinguistic understanding by prompt engineering. Furthermore, many studies have focused on fine-tuning AudioLLMs to elicit their understanding capabilities for specific tasks. Examples include discriminating the human-likeness of audio [51], understanding low-level acoustic qualities [5, 50], and enhancing the assessment of instruction-following in spoken dialogue systems [23, 17]. However, how to improve the ability of AudioLLMs to understand and judge speech naturalness, and how to use the quality assessment capabilities of AudioLLMs as a reward to improve the post-training of speech generation models themselves, remain significantly underexplored.
3 SpeechJudge-Data
Figure 1: SpeechJudge-Data consists of speech pairs (with corresponding text) synthesized by multiple zero-shot TTS models. For each pair, human annotators need to perform (a) a pointwise annotation of text accuracy to assess intelligibility, and (b) a pairwise preference annotation to judge the relative speech naturalness.
Our work is grounded in SpeechJudge-Data, a large-scale human feedback corpus for assessing the intelligibility and naturalness of synthesized speech. Formally, we aim to construct a dataset $\mathcal{D}=\{(t,a_{1},a_{2})\}$, where each triplet comprises a pair of synthesized speech samples $(a_{1},a_{2})$ and the corresponding target text $t$. We instruct annotators to provide pointwise intelligibility and pairwise naturalness preference annotations based on $\mathcal{D}$ (Figure 1).
3.1 Dataset Construction
We employ a diverse set of recent advanced zero-shot TTS models to prepare the dataset $\mathcal{D}$. Formally, for each sample $(t,a_{1},a_{2})$, we denote the synthesized speech $a_{i}$ as being produced by the model $\mathcal{M}_{tts}$, i.e., $a_{i}\sim\mathcal{M}_{tts}(a_{ref},t)$, where $a_{ref}$ is the reference speech.
(a) TTS Models.
(b) Speech References.
(c) Language Settings.
Figure 2: Distribution of SpeechJudge-Data.
Model Selection For $\mathcal{M}_{tts}$, we select the following six models of three architectures to enrich the distribution of the synthetic data (Figure 2(a)): (1) AR-based: ARS [52], CosyVoice2 [14], CosyVoice2-INTP [61], and Ints-INTP [61]. The latter two are released by Zhang et al. [61] as intelligibility-enhanced models. (2) FM-based: F5-TTS. (3) MGM-based: MaskGCT [52].
Prompt Construction To build diverse prompts $(a_{ref},t)$ for TTS, for $a_{ref}$, we adopt both regular and expressive speech samples. The regular samples are randomly selected from the Emilia-Large dataset [19]. The expressive samples are sourced from corpora rich in paralinguistics, including the emotional corpora: ParaSpeechCaps [9], the accented corpora: L2-Arctic [63] and KeSpeech [46], the whisper samples from an in-house corpus, and the character voices from video games Genshin Impact [43]. We display the detailed distribution of speech references in Figure 2(b).
The target text $t$ paired with each $a_{ref}$ is constructed as follows: For regular $a_{ref}$ samples, we randomly sample transcriptions from the Emilia-Large dataset [19]. These are then refined using DeepSeek-V3 [10] to correct typos and normalize punctuations. For expressive $a_{ref}$ samples, we instruct DeepSeek-V3 to generate several scripts in different writing styles, tailored to the topic of $a_{ref}$ (see Appendix 7.1 for more details). The languages of the target texts included Chinese (zh), English (en), and Chinese-English code-switching (mixed). For the combinations $(a_{ref},t)$, we include both monolingual settings (en2en and zh2zh) and cross-lingual settings (zh2en, en2zh, zh2mixed, and en2mixed), where zh2en denotes Chinese $a_{ref}$ with English $t$, and similarly for others. The distribution of the language settings of $(a_{ref},t)$ is shown in Figure 2(c).
[…]
3.2 Human Annotation
Given a sample $(t,a_{1},a_{2})$, human annotators are instructed to perform both pointwise intelligibility and pairwise naturalness annotations (Figure 1). For intelligibility, annotators perform a binary classification to determine whether the speech ($a_{1}$ and $a_{2}$) accurately reads the text $t$ without any content insertion, omission, or mispronunciation. For naturalness, they perform a five-scale Comparative Mean Opinion Score (CMOS) annotation to determine which of the two audio clips ($a_{1}$ or $a_{2}$) sounds more natural and human-like. We recruit human annotators and provide them with training. The detailed annotation guidelines are provided in Appendix 8.
Statistics We recruit 69 annotators and conduct annotations over two months. The resulting constructed dataset $D$, which we denote as SpeechJudge-Data (raw), contains 99K $(t,a_{1},a_{2})$ samples, with each sample receiving an average of 2.49 annotations from different labelers. The market value of this annotation scale is estimated at over 500K RMB (about 70K USD). Based on the raw dataset, we also construct several subsets for analysis and reward model training. We provide detailed descriptions of each subset and its applications in the following sections and in Appendix 7.2.
Human Agreement Analysis We analyze the human annotations for naturalness in this section; discussions regarding intelligibility are provided in Appendix 8.2. For naturalness annotations, we evaluate the inter-annotator agreement across our constructed dataset. To simplify the analysis, given the sample $(t,a_{1},a_{2})$, we transform the five-scale naturalness scale (CMOS) into a ternary classification system: either $a_{1}$ is better, $a_{2}$ is better, or their quality is a Tie. Based on this simplified classification, we categorize the annotation results into four distinct levels of agreement [^1] Note: Each sample of SpeechJudge-Data is independently annotated by a minimum of two and a maximum of three annotators (Appendix 8).: (1) Full Agreement (FA): A consensus is reached among all annotators, with all ratings pointing to the same outcome (e.g., “2A”, “3A”, “2B”, “3B”). We use “2A” to indicate that two annotators both rated $a_{1}$ as better, while “3B” denotes three annotators all rating $a_{2}$ as better. (2) Weak Agreement (WA): This level captures cases where two annotators agree on a specific polarity, while the third annotator marks a Tie (e.g., “2A+1T”, “2B+1T”). We also include the “2T+1A” and “2T+1B” cases in this level. (3) Weak Disagreement (WD): This occurs when two annotators’ ratings share the same polarity, but the third’s rating is the opposite (e.g., “2A+B”, “2B+A”). (4) Full Disagreement (FD): This represents a complete lack of consensus, where all three annotators provide different classifications, denoted as “1A+1B+1T”.
Figure 3: Distribution of SpeechJudge-Data on different levels of human agreement.
In Figure 3, we demonstrate the distribution of these human agreement levels for the SpeechJudge-Data and its two subsets, regular and expressive (which are defined by their speech references). The figure shows that about 70% of the entire dataset falls into the Full Agreement (51.5%) or Weak Agreement (17.2%) levels. Furthermore, we observe that the expressive subset has a lower agreement level than the regular subset, which suggests that human evaluation of expressive speech generation is inherently a more challenging problem. Besides this sample-level agreement analysis, we also analyze the reliability of individual annotators, and we will discuss this in the Appendix 8.1.
4 SpeechJudge-Eval
To evaluate speech naturalness, existing studies typically organize their own listening tests, which often have inconsistent settings across different papers [24, 2, 13, 56, 28]. Alternatively, previous researchers use proxy MOS predictors, such as UTMOS [41], as an objective metric. However, it remains an underexplored problem whether these metrics can accurately judge the naturalness of more advanced speech generation models [52, 14, 7, 61] and align with human preferences. Motivated by this, we construct a benchmark, SpeechJudge-Eval, specifically for the speech naturalness judgment task.
Table 1: Protocols of different models for naturalness judgment.
Type Protocol
Objective
Metrics
WER $\downarrow$, Naturalness $\uparrow$
SIM $\uparrow$, Naturalness $\uparrow$
FAD $\downarrow$, Naturalness $\uparrow$
MOS
Predictors
MOS $\uparrow$, Naturalness $\uparrow$
Deepfake
Detectors
Fake $\downarrow$, Naturalness $\uparrow$
AudioLLMs Score $\uparrow$, Naturalness $\uparrow$
Prompts of AudioLLMs
We are comparing the naturalness of two models’ outputs. The models need to speak the target text accurately and naturally. Target text: {tt}, Output A: $a_{1}$}, Output B: $a_{2}$}. Analyze the two outputs above, and score them with number from 1 to 10. Note: $\circ$ Please evaluate the naturalness of both audio outputs based on the following criteria: Prosody and Intonation, Pacing and Rhythm, Articulation and Clarity, and Overall Naturalness. $\circ$ After conducting a detailed analysis of each criterion, using the following output template to highlight your conclusion: Output A: X, Output B: X.
    We instruct AudioLLMs using two modes of prompt: plain and CoT. The text in blue is only employed during the CoT mode.
Table 2: Accuracy of speech naturalness judgment across different models on SpeechJudge-Eval.
Model Regular Expressive Total
Objective Metrics
WER 59.3 57.0 57.9
SIM 47.5 42.5 44.5
FAD 50.3 47.5 48.6
MOS Predictor
DNSMOS 61.0 55.8 57.9
UTMOS 54.0 53.5 53.7
Content Enjoyment (CE) 69.3 55.2 60.8
Content Usefulness (CU) 61.3 54.7 57.3
Production Complexity (PC) 39.3 48.7 44.9
Production Quality (PQ) 61.3 54.3 57.1
Deepfake Detectors
AASIST 40.5 50.8 46.7
ADV 35.3 40.3 38.3
AudioLLMs (Open-source)
Phi-4-Multimodal 54.8 58.5 57.0
Qwen2.5-Omni-7B 62.0 59.7 60.6
Kimi-Audio-7B-Instruct 65.5 68.0 67.0
Gemma-3n-E4B-it 49.0 47.7 48.2
Voxtral-Mini-3B-2507 60.0 53.3 56.0
MiDashengLM 58.8 63.5 61.6
MiMo-Audio-7B-Instruct 61.3 49.3 54.1
AudioLLMs (Closed-source)
Gemini-2.5-Flash 73.5 66.2 69.1
Gemini-2.5-Pro 73.0 62.2 66.5
GPT-4o mini Audio 56.3 46.7 50.5
GPT-4o Audio 71.5 64.7 67.4
    We use the protocols of Table 1 to establish judgment rules for different models. The results of AudioLLMs here are obtained using the plain prompt of Table 1.
4.1 Task Description
Task Formulation We formulate the naturalness judgment task as a pairwise comparison, specifically a win-or-lose binary classification task: Given a target text $t$ and a corresponding audio pair $(a_{1},a_{2})$, a model needs to determine which audio has better naturalness. This results in a binary choice: either $a_{1}$ is better or $a_{2}$ is better. We use the human answer as the ground truth, and use Accuracy to measure the judgment performance of a model $\mathcal{M}$ on the evaluation set $\mathcal{D}$:
$\textbf{Accuracy}=\frac{1}{|\mathcal{D}|}\sum_{d=0}^{|\mathcal{D}|}\mathbb{I}(y_{\mathcal{M}}=y_{\mathcal{H}}),$ (1)
where $|\mathcal{D}|$ is the total number of samples in the evaluation set, $y_{\mathcal{M}}$ and $y_{\mathcal{H}}$ represent the answers of the model $\mathcal{M}$ and human for the sample $d$, respectively. $\mathbb{I}$ is the indicator function.
Evaluation Data We sample a subset of the SpeechJudge-Data to create the evaluation set for SpeechJudge-Eval. Specifically, we first select a subset that contains only preference data (i.e., we filter out samples with the “Tie” annotation), and then choose only those with full-agreement-level (FA) samples to ensure a high-quality ground truth. We perform sampling from both the regular and expressive subsets of SpeechJudge-Data and proportionally cover the three target text languages (zh, en, and mixed) within each subset. The final SpeechJudge-Eval dataset consists of 1,000 samples. The construction details of SpeechJudge-Eval and its distribution can be found in Appendix 7.2.
4.2 Benchmark for Different Models
We test the naturalness judgment capability of various models based on SpeechJudge-Eval. We consider four different categories of models, whose evaluation protocols are shown in Table 1:
  1. 1.
    Objective metrics, such as WER [38, 16], SIM [6], and FAD [27] in audio generation tasks. We assume that a better value of these metrics (e.g., lower for WER and FAD; higher for SIM) indicates better naturalness.
  2. 2.
    MOS Predictors, including DNSMOS [40], UTMOS [41], and predictors from audiobox-aesthetics (CE, CU, PC, and PQ) [48]. We assume that a higher MOS score corresponds to better naturalness.
  3. 3.
    Deepfake detectors, which are typically pre-trained on a binary classification task to predict whether an audio is fake or not [25, 49]. We assume that an audio with a lower fake probability should have better naturalness.
  4. 4.
    AudioLLMs, which are employed to test their speech naturalness understanding capabilities in a zero-shot manner [^2] We assume that the adopted AudioLLMs have not been directly trained on the speech naturalness judgment task. Their performance on this benchmark is therefore considered a zero-shot capability.. We include the open-source Phi-4-Multimodal [1], Qwen2.5-Omni [56], Kimi-Audio [28], Gemma-3n [26], Voxtral [32], MiDashengLM [12], Mimo-Audio [53], and the closed-source Gemini-2.5 [8] and GPT-4o [22]. We use the plain prompt of Table 1 to instruct the model to pairwise score the naturalness of two audios. We use their grading to determine the naturalness preference.
The performance of different models on SpeechJudge-Eval is presented in Table 2. A key observation is that speech naturalness judgment is a highly challenging task. The leading model, Gemini-2.5-Flash, still only achieves less than 70% agreement with human preferences. When comparing different models, we find that: (1) common objective metrics and MOS predictors show only a weak correlation with human preferences, often achieving less than 60% accuracy and sometimes performing at the level of a random guess (around 50%). (2) While deepfake detectors are highly effective at distinguishing between machine-generated and human-recorded speech [49, 25], their ability to do so is not well-aligned with the naturalness objective when comparing two generated samples. (3) AudioLLMs demonstrate significant potential for this task. While some models, such as Gemma-3n and GPT-4o mini Audio, perform at a chance level, a number of others achieve an accuracy exceeding 60%. This promising performance motivates us to further leverage these AudioLLMs for the design of a reward model for speech naturalness.
5 SpeechJudge-GRM
Figure 4: SpeechJudge-GRM: (a) We employ Gemini-2.5-Flash as a teacher model to generate CoT rationales for SpeechJudge-Data. We use the samples where Gemini-2.5-Flash’s preference aligns with human as the SFT dataset, while the remaining samples are reserved for the RL stage. (b) We treat the human preference as a verifiable reward to train the GRM with GRPO.
Based on the proposed SpeechJudge-Data, we further explore how to train a reward model capable of accurately capturing human preferences. Specifically, we propose SpeechJudge-GRM, where we leverage the inherent audio understanding capabilities of AudioLLMs (specifically, Qwen2.5-Omni-7B [56]) to elicit their speech naturalness judgment capability. Compared to the classic BTRM [4], the key strengths of GRM are its ability to enable Chain-of-Thought (CoT) reasoning and its support for test-time computation via majority voting, which ultimately leads to improved preference judgment performance [59].
5.1 Methodology
We develop SpeechJudge-GRM based on Qwen2.5-Omni-7B (Thinker) [56]. Inspired by the powerful capabilities of RL with the verifiable reward (RLVR) [42, 11], our natural initial approach is to treat the human preference $y_{\mathcal{H}}$ for the pair $(a_{1},a_{2})$ as a verifiable reward, and launch a RLVR training based on Qwen2.5-Omni. However, in practice, we find that the instruction-following reasoning capabilities of Qwen2.5-Omni are very weak (more detailed discussions can be found in Appendix 10). Therefore, we adopt a two-stage post-training process (“SFT + RL”) to develop SpeechJudge-GRM (Figure 4). We describe the details as follows.
Table 3: Accuracy of speech naturalness judgment of SpeechJudge-GRM.
Model Regular Expressive Total
Qwen2.5-Omni-7B 62.0 59.7 60.6
Gemini-2.5-Flash 73.5 66.2 69.1
SpeechJudge-BTRM 77.5 69.5 72.7
SpeechJudge-GRM (SFT) 77.8 73.7 75.3
w/ Voting@10 77.4 77.6 77.6
SpeechJudge-GRM (SFT+RL) 79.0 76.0 77.2
w/ Voting@10 80.5 78.7 79.4
    Our evaluation is conducted on SpeechJudge-Eval (like Table 2). w/ Voting@10: For each prompt, the GRM generates 10 outputs, and we use the majority voting from these 10 outputs as the final result.
Figure 5: Subjective evaluation of using SpeechJudge-GRM for high-naturalness sample selection. Human subjects compare a best-of-100 output of Qwen2.5-Omni-7B (Talker), chosen by either SpeechJudge-BTRM or SpeechJudge-GRM, against a randomly output.
SFT Stage We consider SFT as a “cold start” stage to improve the Qwen2.5-Omni’s instruction-following, reasoning, and speech naturalness understanding capabilities. We select Gemini-2.5-Flash [8]—one of the leading closed-source models on SpeechJudge-Eval (Table 2)—to serve as a teacher model, and instruct it to generate the CoT data. Specifically, for each sample $d=(t,a_{1},a_{2},y_{\mathcal{H}})$ from SpeechJudge-Data, we use the CoT prompt from Table 1 (denoted as $\mathbf{I}_{CoT}$) to instruct Gemini-2.5-Flash to generate a rationale-based output (denoted as $\mathbf{O}_{teacher}$). We then extract the preference judgment ($y_{\mathcal{M}}$) from this output. For samples where Gemini-2.5-Flash’s preference is consistent with the human (i.e., $y_{\mathcal{M}}=y_{\mathcal{H}}$), we concatenate the CoT prompt and the model’s output, $\mathbf{I}_{CoT},\mathbf{O}_{teacher}]$, to create a data point for our SFT dataset. Conversely, we consider the sample $d$ a challenging case and reserve the prompt $\mathbf{I}_{CoT}$ for the second-stage RL dataset. During the SFT stage, for each training sample $\mathbf{I}_{CoT},\mathbf{O}_{teacher}]$, we perform the next token prediction only on the segment $\mathbf{O}_{teacher}$.
RL Stage We treat the annotated human preference as a verifiable reward, and, building on the SFT model, we further trained it using the GRPO algorithm [42]. Specifically, for each sample $d=(t,a_{1},a_{2},y_{\mathcal{H}})$ in the RL dataset, we adopt the CoT prompt to instruct the policy model to conduct multiple rollouts during each iteration. For the $i$-th rollout, we parse the model’s preference for $(a_{1},a_{2})$, denoted as $y_{\mathcal{M}}^{i}$. Following [34], we use an accuracy-based rule to calculate the reward: the reward is 1 if $y_{\mathcal{M}}^{i}=y_{\mathcal{H}}$, and -1 otherwise. In other words, during the RL stage, we only constrain the model’s final naturalness judgment to align with human preferences, allowing the model to autonomously optimize its reasoning and rationale generation capabilities.
We denote the training dataset of SpeechJudge-GRM as SpeechJudge-Data (train). Its construction process is as follows (see Appendix 7.2 for more details). Based on the raw SpeechJudge-Data, we first filter out all samples at the Full Disagreement (FD) level. For the other samples—at the FA, WA, and WD levels—we apply a majority voting principle among annotators to determine the final label for each. We then further exclude samples with a “Tie” label, using only the remaining preference data to form the SpeechJudge-Data (train). We use LoRA [20] to fine-tune the GRM during both the SFT and RL stages. Other experimental setup details are provided in Appendix 11.
5.2 Effectiveness of SpeechJudge-GRM on Naturalness Judgement
To verify the effectiveness of SpeechJudge-GRM for naturalness judgment, we evaluate it on the SpeechJudge-Eval benchmark. We develop SpeechJudge-BTRM as a baseline, which utilizes the BTRM paradigm [4, 39] by adding a linear layer on Qwen2.5-Omni-7B (Thinker) to produce a single scalar reward prediction. SpeechJudge-BTRM also uses LoRA fine-tuning and uses the same training data as SpeechJudge-GRM.
From the results of Table 3, we can observe that: (1) The SpeechJudge-BTRM achieves a 72.7% agreement with human preferences on SpeechJudge-Eval, a level of performance comparable to the initial development of BTRMs in the textual LLM RLHF field [44, 3, 37]. (2) After conducting SFT training with the CoT data, the accuracy of SpeechJudge-GRM (SFT) reaches 75.3%. Besides, further RLVR training improves the final model SpeechJudge-GRM (SFT+RL) to an accuracy of 77.2%. (3) Due to the generative nature of the GRM, we can further enhance the accuracy of SpeechJudge-GRM using inference-time scaling. For example, by using majority voting across 10 outputs instead of just one, the accuracy is improved by approximately 2 percentage points (75.3% $\rightarrow$ 77.6%; 77.2% $\rightarrow$ 79.4%). These results collectively verify the effectiveness of our proposed SpeechJudge-GRM for judging speech naturalness.
5.3 High-Quality Sample Selection based on SpeechJudge-GRM
We investigate the effect of SpeechJudge-based reward models for high-quality sample selection. We use the hard cases from SeedTTS-Eval [2] and the code-switching cases from Amphion-TTS-Eval [60] as target texts. For each text, we instruct the Qwen2.5-Omni-7B (Talker) [57] to generate 100 speeches. We then ask human subjects to compare the best-of-100 output—as selected by either SpeechJudge-BTRM or SpeechJudge-GRM—against a randomly sampled output. The evaluation measures the win/lose/tie ratios based on speech naturalness. From Figure 5, we observe that the best-of-100 samples selected by both SpeechJudge-BTRM and SpeechJudge-GRM are more likely to outperform a randomly selected sample from the same set. This finding demonstrates the advantage of using the SpeechJudge-Data corpus for training human-aligned reward model. Furthermore, SpeechJudge-GRM exhibits better performance than SpeechJudge-BTRM, which highlights the superiority of the proposed GRM.
5.4 Post-Training of Zero-Shot TTS based on SpeechJudge-GRM
Model T-ACC N-CMOS
Qwen2.5-0.5B-TTS 84.0% 0.00 $_{\scriptscriptstyle\pm\text{0.00}}$
w/ INTP 87.0% 0.18 $_{\scriptscriptstyle\pm\text{0.07}}$
w/ SpeechJudge-Data 91.0% 0.16 $_{\scriptscriptstyle\pm\text{0.08}}$
w/ SpeechJudge-GRM (offline) 91.0% 0.21 $_{\scriptscriptstyle\pm\text{0.12}}$
w/ SpeechJudge-GRM (online) 90.0% 0.25 $_{\scriptscriptstyle\pm\text{0.09}}$
(a) Text Accuracy (T-ACC) and Naturalness CMOS (N-CMOS).
(b) Win/Lose/Tie of speaker similarity after post-training.
Figure 6: Post-training of Qwen2.5-0.5B-TTS based on SpeechJudge. We display the objective results (WER and SIM) in Table 6 of Appendix 12.2.
We investigate the effect of using SpeechJudge-GRM as a reward function for post-training of TTS model. Specifically, we develop a new zero-shot TTS model, Qwen2.5-0.5B-TTS, to serve as the base model, which was not involved in the construction of the SpeechJudge-Data. This model is based on Qwen2.5-0.5B [57], adopts the classic two-stage “AR+Diffusion” architecture [2, 13], uses the speech tokenizer from DualCodec [31], and is pre-trained on the Emilia dataset [19].
Based on this pre-trained model, we design four comparative methods: (1) w/ INTP: We use the intelligibility preference dataset, INTP [61], to perform offline DPO alignment [39]. (2) w/ SpeechJudge-Data: We use the SpeechJudge-Data (train) to perform offline DPO alignment. (3) w/ SpeechJudge-GRM (offline): We use SpeechJudge-GRM as an offline preference data annotator. We take all speech pairs from the INTP dataset and re-annotate their preference labels using SpeechJudge-GRM, then perform offline DPO alignment on the resulting data. (4) w/ SpeechJudge-GRM (online): We use SpeechJudge-GRM as a reward function for the online DPO algorithm [18]. The training data consists of only the prompts from INTP (i.e., the target texts and speech references for zero-shot TTS).
We use SeedTTS-Eval [2] and Amphion-TTS-Eval [61, 60, 62] as evaluation sets. We present the objective results (WER and SIM) in Table 6 and the subjective results in Figure 6. We observe that both intelligibility and naturalness are enhanced for all the four methods after post-training. Additionally, the post-training method based on SpeechJudge-GRM achieves a greater improvement in naturalness (Figure 6(a)). Besides, the SpeechJudge-based methods could match or lead to a slight improvement in speaker similarity (Figure 6(b)).
6 Conclusion
In this work, we address the challenge of aligning speech synthesis with human perception of naturalness by introducing SpeechJudge, a suite including a large-scale dataset (SpeechJudge-Data), a challenging benchmark (SpeechJudge-Eval), and a generative reward model (SpeechJudge-GRM). Our benchmark reveals that even top AudioLLMs struggle with naturalness judgment, achieving less than 70% human agreement. In contrast, our proposed SpeechJudge-GRM reaches 77.2% accuracy (up to 79.4% with inference-time scaling @10), outperforming the classic Bradley-Terry reward models (72.7%). We also demonstrate its practical utility as a reward function to effectively enhance TTS model naturalness via post-training. While our primary focus is on naturalness, future work could extend this framework to other subjective attributes like speaker similarity and emotional expressiveness, potentially through multi-objective reward modeling. By releasing our resources, we aim to catalyze research in building more human-aligned speech generation systems.